# Uncertainty in Graph Neural Networks: A Survey

### v1

@@ 6 Future directions
Identifying fine-grained graph uncertainty. While current research on graph data has asserted their ability to differentiate between various types and sources of uncertainty (Section 2), only a few efforts have been made to identify and quantify uncertainty tailored specifically for graph components. For example, (bojchevski2017deep, ; liu2022ud, ) have assumed that heterophily, the conflict between two connected nodes, is correlated with uncertainty in node classification. These attempts demonstrate that fine-grained graph uncertainty is essential for downstream tasks. Therefore, it is promising to develop other unexplored fine-grained uncertainties that contribute to such specific tasks. For instance, researchers might seek to identify specific types of graph distribution shifts, such as shifts in graph node features, labels, and graph structure. Decomposing distributional uncertainty at a more granular level could aid in distinguishing and leveraging the corresponding shifts, thereby enhancing OOD detection and generalization. Furthermore, different forms of finer-grained graph uncertainty should be comparable in scale. By comparing their corresponding uncertainty measures, we can determine whether the distributional uncertainty arising from node feature or graph topology, or the Gnn modeling is dominant. This contributes to uncertainty estimation with improved explainability by indicating how each component in graph data, Gnn modeling, and inference process contributes to the final Gnn predictive uncertainty.
@@
Constructing ground-truth datasets and unified evaluation metrics. The critical unresolved issue of the lack of ground-truth datasets and metrics for evaluation, as discussed in Section 3.4, persists. While aleatoric and distributional uncertainty can be artificially introduced into synthetic graph data with predefined underlying distributions or generating processes, synthetic data has difficulty in simulating complex real-world scenarios, with diverse and disentangled sources of uncertainty. Besides, it is also challenging to obtain ground-truth values for model uncertainty, as it depends on the selection of models. On the other hand, though humans could be involved in evaluating ground-truth aleatoric uncertainty for real-world data, biases may still arise due to varying levels of human cognitive abilities. Additionally, the lack of unified evaluation metrics for different tasks, such as classification or regression tasks, further complicates the evaluation process. Thus, uncertainty across different tasks is not directly comparable. However, as a promising direction, it may be feasible to assess whether uncertainty estimation enhances well-defined metrics for the final task rather than directly evaluating the quality of uncertainty estimation itself.
@@
Quantifying task-oriented uncertainty. The existing literature lack a systematic investigation aimed at identifying the most effective uncertainty method for downstream tasks introduced in Section 4. This is partly attributable to the fact that researchers in a specific domain often stick to a specific way to quantify uncertainty. For instance, in graph active learning, entropy is frequently used for uncertainty estimation without additional elucidation. It remains unclear whether alternative uncertainty measures, such as mutual information, either independently or in combination with other selection criteria, would yield superior performance. Another interesting direction is whether the separation of aleatoric and epistemic uncertainty is necessary. For instance, in other active learning tasks, epistemic uncertainty measures generally outperform total uncertainty ones, while aleatoric uncertainty generally excels in failure detection (kahl2024values, ). However, in graph active learning and outlier detection, the common measures capture predictive uncertainty instead of separating it. Therefore, there is a need for systematic validation of various uncertainty types, quantification methods, and evaluation techniques across a wide range of pertinent downstream tasks.
@@
Designing better methods for real-world applications. In Section 3, we have introduced multiple categories of uncertainty quantification methods and showed their strengths and weaknesses. In real-world applications, several aspects need to be taken into primary consideration. First of all, whether methods can effectively capture specific sources and types of uncertainty that are relevant to particular types of tasks. For example, Dirichlet-based methods directly model distributional uncertainty, thus showing superior performance over other methods in graph OOD detection. Secondly, computational efficiency and difficulty of application to existing prediction frameworks are also of vital importance. For instance, though ensemble models outperform most other methods in both model prediction and uncertainty estimation, they are rarely applied to real-world problems due to their high computational costs. One potential direction is to explore graph ensemble distillation (zhang2020reliable, ) as a cost-efficient alternative, which uses smaller Gnns trained on sub-problems and teaches one student Gnn to handle OOD and estimate uncertainty (hinton2015distilling, ). Lastly, one should explore whether the proposed methods can consistently achieve reliable performance across diverse settings. For example, the maximal softmax probability is widely applied in real-world problems due to time efficiency but suffers from under-confidence and is subject to the effect of distributional shifts. In conclusion, it is an exciting direction to develop more efficient, easy-to-apply, robust graph uncertainty quantification methods that can identify diverse fine-grained sources and types of uncertainty.

### v2

@@ 6 Future directions
Identifying fine-grained graph uncertainty. While current research on graph data has asserted their ability to differentiate between various types and sources of uncertainty (Section 2), only a few efforts have been made to identify and quantify uncertainty tailored specifically for graph components. For example, (Bojchevski & Günnemann, 2017; Liu et al., 2022e) have assumed that heterophily, the conflict between two connected nodes, is correlated with uncertainty in node classification. These attempts demonstrate that fine-grained graph uncertainty is essential for downstream tasks. Therefore, it is promising to develop other unexplored fine-grained uncertainties that contribute to such specific tasks. For instance, researchers might seek to identify specific types of graph distribution shifts, such as shifts in graph node features, labels, and graph structure. Decomposing distributional uncertainty at a more granular level could aid in distinguishing and leveraging the corresponding shifts, thereby enhancing OOD detection and generalization. Furthermore, different forms of finer-grained graph uncertainty should be comparable in scale. By comparing their corresponding uncertainty measures, we can determine whether the distributional uncertainty arising from node feature or graph topology, or the Gnn modeling is dominant. This contributes to uncertainty estimation with improved explainability by indicating how each component in graph data, Gnn modeling, and inference process contributes to the final Gnn predictive uncertainty.
@@
Constructing ground-truth datasets and unified evaluation metrics. The critical unresolved issue of the lack of ground-truth datasets and metrics for evaluation, as discussed in Section 3.4, persists. While aleatoric and distributional uncertainty can be artificially introduced into synthetic graph data with predefined underlying distributions or generating processes, synthetic data has difficulty in simulating complex real-world scenarios, with diverse and disentangled sources of uncertainty. Besides, it is also challenging to obtain ground-truth values for model uncertainty, as it depends on the selection of models. On the other hand, though humans could be involved in evaluating ground-truth aleatoric uncertainty for real-world data, biases may still arise due to varying levels of human cognitive abilities. Additionally, the lack of unified evaluation metrics for different tasks, such as classification or regression tasks, further complicates the evaluation process. Thus, uncertainty across different tasks is not directly comparable.
Currently, a feasible alternative is to assess whether uncertainty estimation enhances well-defined metrics for the final task performance rather than directly evaluating the quality of uncertainty estimation. However, there is a gap in understanding whether performance enhancements are solely due to the selection of quantification methods or if they ideally result from an appropriate match between quantification methods and corresponding uncertainty types, where the uncertainty type itself aids the task. Bridging this gap still requires further efforts.
@@
Quantifying task-oriented uncertainty. The existing literature lack a systematic investigation aimed at identifying the most effective uncertainty method for downstream tasks introduced in Section 4. This is partly attributable to the observation that researchers in a specific domain tend to follow a specific method to quantify uncertainty. For instance, in graph active learning, entropy is frequently used for uncertainty estimation without additional elucidation. It remains unclear whether alternative uncertainty measures, such as mutual information, either independently or in combination with other selection criteria, would yield superior performance. Another interesting direction is whether the separation of aleatoric and epistemic uncertainty is necessary. For instance, in other active learning tasks, epistemic uncertainty measures generally outperform total uncertainty ones, while aleatoric uncertainty generally excels in failure detection (Kahl et al., 2024). However, in graph active learning and outlier detection, the common measures capture predictive uncertainty instead of separating it. Therefore, there is a need for systematic validation of various uncertainty types, quantification methods, and evaluation techniques across a wide range of pertinent downstream tasks.
@@
Designing better methods for real-world applications. In Section 3, we have introduced multiple categories of uncertainty quantification methods and showed their strengths and weaknesses. In real-world applications, several aspects need to be taken into primary consideration. First of all, whether methods can effectively capture specific sources and types of uncertainty that are relevant to particular types of tasks. For example, Dirichlet-based methods directly model distributional uncertainty, thus showing superior performance over other methods in graph OOD detection. Secondly, computational efficiency and difficulty of application to existing prediction frameworks are also of vital importance. For instance, though ensemble models outperform most other methods in both model prediction and uncertainty estimation, they are rarely applied to real-world problems due to their high computational costs. One potential direction is to explore graph ensemble distillation (Zhang et al., 2020) as a cost-efficient alternative, which uses smaller Gnns trained on sub-problems and teaches one student Gnn to handle OOD and estimate uncertainty (Hinton et al., 2015). Lastly, one should explore whether the proposed methods can consistently achieve reliable performance across diverse settings. For example, the maximal softmax probability is widely applied in real-world problems due to time efficiency but suffers from under-confidence and is subject to the effect of distributional shifts. In conclusion, it is an exciting direction to develop more efficient, easy-to-apply, robust graph uncertainty quantification methods that can identify diverse fine-grained sources and types of uncertainty.
@@ 7 Conclusions
Acknowledgments
This work is supported in part by NSF under grants III-2106758, and POSE-2346158.